# Eisenstein degeneration of Euler systems

### v1

@@ 2.2. Families of representations
(1) $0\longrightarrow\mathbf{Q}_{p}(\tau\varepsilon^{-1-r})\longrightarrow V^{c}(f_{\beta})\longrightarrow\mathbf{Q}_{p}(\psi)\longrightarrow 0,$
and from here we get another short exact sequence of $L$-linear Galois representations for $V(f_{\beta})$.
(2) $0\longrightarrow\mathbf{Q}_{p}(\psi)\longrightarrow V(f_{\beta})\longrightarrow\mathbf{Q}_{p}(\tau\varepsilon^{-1-r})\longrightarrow 0$
@@
(3) $0\longrightarrow\mathbf{Q}_{p}(\tau^{-1}\varepsilon^{1+r})\longrightarrow V(f_{\beta})^{*}\longrightarrow\mathbf{Q}_{p}(\psi^{-1})\longrightarrow 0,$
@@
(4) $0\longrightarrow\mathbf{Q}_{p}(\psi^{-1})\longrightarrow V^{c}(f_{\beta})^{*}\longrightarrow\mathbf{Q}_{p}(\tau^{-1}\varepsilon^{1+r})\longrightarrow 0$
There is a natural map $V^{c}(f_{\beta})^{*}\to V(f_{\beta})^{*}$ whose image is the $\tau^{-1}\varepsilon^{1+r}$ subrepresentation of the latter.
@@
The Coleman–Mazur–Buzzard cuspidal eigencurve $\mathcal{C}^{0}$ of tame level $N$ is smooth at $f_{\beta}$ and locally étale over the weight space.
This assumption is automatic when $f$ is decent and non-critical; as discussed in [LRS24, §A], these are fairly general conditions. Moreover, this means that may take a closed disc $U$ around $f_{\beta}$ mapping isomorphically to its image in the weight space.
Let ${\mathbf{f}}$ be the universal eigenform over $U$, which is an overconvergent cuspidal eigenform with coefficients in $\mathcal{O}(U)$. We also write $\beta_{{\mathbf{f}}}\in\mathcal{O}(U)^{\times}$ for the Hecke $U(p)$-eigenvalue of ${\mathbf{f}}$. There is also a canonical Galois pseudo-character $t_{{\mathbf{f}}}\colon G_{\mathbf{Q}}\rightarrow\mathcal{O}(U)$ satisfying
@@
Then, we have two $\mathcal{O}(U)$-linear Galois representations which both have traces equal to $t_{{\mathbf{f}}}$, and two more whose traces are the dual $t_{\mathbf{f}}^{*}$. As shown in [LZ16, §4], there exist sheaves $\mathscr{D}_{U}$ and $\mathscr{D}^{\prime}_{U}$ of $\mathcal{O}(U)$-modules on the modular curve $Y$, which allow us to produce the corresponding Galois representations. We refer the reader to loc. cit. for the precise definitions of the sheaves.
Definition 2.4.
We define
[…]
and $V({\mathbf{f}})$ carries a $\mathcal{O}(U)$-linear Galois representation whose trace is $t_{{\mathbf{f}}}$. By [BC06, Thm. 1], $t_{{\mathbf{f}}}$ has maximal reducibility ideal. Then, it follows that the fiber of $V({\mathbf{f}})$ at $k=r$, which is exactly $V(f_{\beta})$, must be a non-split extension. The same applies to the other three modules, showing that all four of $V(f_{\beta}),V^{c}(f_{\beta}),V(f_{\beta})^{*}$ and $V^{c}(f_{\beta})^{*}$ are non-split extensions in the setting of Soulé’s theorem.
@@ 2.3. Crystalline periods
We begin by establishing some notation. Fix an embedding $\overline{\mathbf{Q}}\hookrightarrow\overline{\mathbf{Q}}_{p}$, that allow us to see $\operatorname{Gal}(\overline{\mathbf{Q}}_{p}/\mathbf{Q}_{p})$ as a subgroup of $\operatorname{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$. We write $\mathcal{R}$ for the Robba ring over $\mathbf{Q}_{p}$, and let $\mathcal{R}_{U}=\mathcal{R}\mathop{\hat{\otimes}}\mathcal{O}_{U}$. A $(\varphi,\Gamma)$-module is endowed with a Frobenius action (denoted by $\varphi$) and a Galois action, for which we take a topological generator $\gamma$ of $\operatorname{Gal}(\overline{\mathbf{Q}}_{p}/\mathbf{Q}_{p})$. As usual, let $t\in\mathcal{R}$ be the period for the cyclotomic character, so $\varphi(t)=pt$ and $\gamma(t)=\varepsilon(\gamma)t$. For $\delta\in\mathcal{O}(U)^{\times}$, let $\mathcal{R}_{U}(\delta)$ be the rank 1 $(\varphi,\Gamma)$-module over $\mathcal{R}_{U}$ generated by an element $e$ which is $\Gamma$-invariant and satisfies $\varphi(e)=\delta e$. Finally, when $D$ is a $(\varphi,\Gamma)$-module over $\mathcal{R}$ (or $\mathcal{R}_{U}$), write $\mathbf{D}_{\mathrm{cris}}(D)=D[1/t]^{\Gamma}$, with its filtration $\operatorname{Fil}^{n}\mathbf{D}_{\mathrm{cris}}(D)=(t^{n}D)^{\Gamma}$. Let $D({\mathbf{f}})^{*}=\mathbf{D}^{\dagger}_{\mathrm{rig}}(V({\mathbf{f}})^{*})$, and similarly $D^{c}({\mathbf{f}})^{*}$.
@@
is a finitely-generated $\mathcal{O}(U)$-module, by the main theorem of [KPX14]. In particular, it is free of rank 1, and there is a Zariski-dense set of $x\in\mathcal{O}(U)$ where the fiber is 1-dimensional. Let $\mathcal{F}^{+}D({\mathbf{f}})^{*}$ be the image of a generator of this map, and $\mathcal{F}^{-}$ the quotient, so that we have a short exact sequence
$0\to\mathcal{F}^{+}D({\mathbf{f}})^{*}\to D({\mathbf{f}})^{*}\to\mathcal{F}^{-}D({\mathbf{f}})^{*}\to 0$
If $f_{\beta}$ is non-critical, the submodule $\mathcal{F}^{+}D(f_{\beta})^{*}$ is saturated, and $\mathcal{F}^{-}D({\mathbf{f}})^{*}$ is free of rank 1 over $\mathcal{R}_{U}$. Further, ±$\mathcal{F}^{\pm}D({\mathbf{f}})^{*}$ define a triangulation of $D({\mathbf{f}})^{*}$ over $U$.
@@
$b_{{\mathbf{f}}}^{+}:\mathbf{D}_{\mathrm{cris}}\left(\mathcal{F}^{+}D({\mathbf{f}})^{*}(-1-{\mathbf{k}})\right)\cong\mathcal{O}_{U}.$
@@
and we have a duality between $V(f_{\beta})^{*}_{\mathrm{quo}}$ and $V^{c}(f_{\beta})_{\mathrm{quo}}$; so we should regard $b_{f_{\beta}}^{+}$ as a basis of the space
$\mathbf{D}_{\mathrm{cris}}(V^{c}(f_{\beta})_{\mathrm{quo}})\cong H^{1}_{\mathrm{dR},\mathrm{c}}(Y,\mathscr{V}_{r})[T=f_{\beta}]=H^{1}(X,\omega^{-r}(-\mathrm{cusps}))[T=f_{\beta}].$
Hence, there exists some scalar $c_{f_{\beta}}\in L^{\times}$ such that
@@
but a priori we do not have any control over that constant.
@@
$a_{{\mathbf{f}}}^{-}:\mathbf{D}_{\mathrm{cris}}\left(\mathcal{F}^{-}D({\mathbf{f}})^{*}\right)\cong\mathcal{O}_{U}.$
If $g$ is a non-critical classical specialization of ${\mathbf{f}}$, then $a_{{\mathbf{f}}}^{-}$ specializes to $d_{g}\eta_{g}$, where $d_{g}$ is some arbitrary non-zero constant. At $X=0$, we have an isomorphism
$\mathbf{D}_{\mathrm{cris}}(\mathcal{F}^{-}D(f_{\beta})^{*}(-1-r))\cong\mathbf{D}_{\mathrm{cris}}(V(f_{\beta})^{*}_{\mathrm{sub}}),$
so we should regard $a_{f_{\beta}}^{-}$ as an object playing the role of $\omega_{f_{\beta}}$.
@@ 3. Circular units
As in the introduction, let $\mathcal{H}_{\Gamma}$ be the distribution algebra of $\Gamma\cong\mathbf{Z}_{p}^{\times}$ (with $L$-coefficients), and, for our further use, let ${\mathbf{j}}\colon\Gamma\hookrightarrow\mathcal{H}_{\Gamma}^{\times}$ be the universal character, which we regard as a character of the Galois group by composition with the cyclotomic character. The Kubota–Leopoldt $p$-adic $L$-function attached to $\chi$, denoted by $L_{p}(\chi)$, is an element of $\mathcal{H}_{\Gamma}$ which interpolates special values of the Dirichlet $L$-function of $\chi$. We refer [Das16, §3] for a more detailed treatment of the properties of this $p$-adic $L$-function with our current conventions, which slightly differ from [Col00].
@@
The element $L_{p}(\chi)$ belongs to the Iwasawa algebra $\mathcal{O}_{\chi,p}[\![\mathbb{Z}_{p}^{\times}]\!]$, which is a product of $(p-1)$ subrings, indexed by $\mathbb{Z}/(p-1)\mathbb{Z}$, each of them isomorphic to $\mathcal{O}_{\chi,p}[\![T]\!]$.
Then, for any $t\in\mathbb{Z}/(p-1)\mathbb{Z}$ with $(-1)^{t}=\chi(-1)$, there exists a continuous function $L_{p,t}(\chi)\colon\mathbb{Z}_{p}\rightarrow\mathbb{C}_{p}$ such that, for all integers $k\geq 0$ with $k\equiv t\pmod{p-1}$, we have $L_{p,t}(\chi,k)=(1-\chi(p)p^{k-1})\cdot L(\chi,1-k)$. For the subrings corresponding to the sign condition $(-1)^{t}=-\chi(-1)$, we define the corresponding $L_{p,t}(\chi)$ using the interpolation property along the positive integers, thus having $(p-1)$ different $p$-adic $L$-function which can be packaged in a single element $L_{p}(\chi)$ of $\mathcal{O}_{\chi,p}[\![\mathbb{Z}_{p}^{\times}]\!]$.
With these conventions, the Kubota–Leopoldt $p$-adic $L$-function satisfies a functional equation, which is discussed in [Das16, §3.2]. However, although the Kubota–Leopoldt $p$-adic $L$-function may be defined along any of the $(p-1)$ discs of the weight space, regardless of any parity constraint, the cohomology classes can only be defined over half of them.
@@
Here, as usual, we identify $\chi$ with a character of the Galois group. It is characterized by the property that for any integer $j$, we recover $V(\chi)^{*}(j)=\mathcal{O}_{\chi,p}(\chi^{-1})(-j)$ by specializing at $j$ (viewed as a ring homomorphism $\mathcal{H}(\Gamma)\to\mathbf{Q}_{p}$).

### v3

@@ 2.2. Families of representations
$0\longrightarrow L(\tau\varepsilon^{-1-r})\longrightarrow V^{c}(f_{\beta})\longrightarrow L(\psi)\longrightarrow 0,$
and from here we get another short exact sequence of $L$-linear Galois representations for $V(f_{\beta})$:
$0\longrightarrow L(\psi)\longrightarrow V(f_{\beta})\longrightarrow L(\tau\varepsilon^{-1-r})\longrightarrow 0.$
@@
$0\longrightarrow L(\tau^{-1}\varepsilon^{1+r})\longrightarrow V(f_{\beta})^{*}\longrightarrow L(\psi^{-1})\longrightarrow 0,$
@@
$0\longrightarrow L(\psi^{-1})\longrightarrow V^{c}(f_{\beta})^{*}\longrightarrow L(\tau^{-1}\varepsilon^{1+r})\longrightarrow 0.$
There is a natural map $V^{c}(f_{\beta})^{*}\to V(f_{\beta})^{*}$ whose image is the $\tau^{-1}\varepsilon^{1+r}$ subrepresentation of the latter; see e.g. the discussion around [BD15, Rk. 5.10].
@@
The Coleman–Mazur–Buzzard cuspidal eigencurve $\mathcal{C}^{0}$ of tame level $\Gamma_{1}(N)$ is smooth at $f_{\beta}$ and locally étale over the weight space.
This assumption is automatic when $f$ is decent and non-critical; as discussed in [LRS24, §A], these are fairly general conditions. Moreover, this means that may take an affinoid open $U=\mathrm{Spa}(A_{U},A_{U}^{\circ})$ around $f_{\beta}$ mapping isomorphically to its image in the weight space.
Let $\mathbf{f}$} be the universal eigenform over $U$, which is an overconvergent cuspidal eigenform with coefficients in $A_{U}$. We also write $\beta_{{\mathbf{f}}}\in A_{U}^{\times}$ for the Hecke $U(p)$-eigenvalue of $\mathbf{f}$}. There is also a canonical Galois pseudo-character $t_{{\mathbf{f}}}\colon G_{\mathbf{Q}}\rightarrow A_{U}$ satisfying
@@
According to Proposition A.11 and Remark A.13, there exist Galois representations $V({\mathbf{f}})$, $V^{c}({\mathbf{f}})$ (resp., $V({\mathbf{f}})^{*}$, $V^{c}({\mathbf{f}})^{*}$), free of rank-$2$ over $A_{U}$ whose traces equal to $t_{{\mathbf{f}}}$ (resp., $t_{{\mathbf{f}}}^{*}$). Moreover, by Corollary A.12, we have injections of Galois representations
$V^{c}({\mathbf{f}})\hookrightarrow V({\mathbf{f}})\quad(\text{resp., }V^{c}({\mathbf{f}})^{*}\hookrightarrow V({\mathbf{f}})^{*}).$
We refer the reader to Appendix A for the detailed construction of these representations.
Let $Y$ be the modular curve of level $\Gamma_{1}(N)\cap\Gamma_{0}(p)$ over $\mathbf{Q}$ and let $\overline{Y}$ be the base change of $Y$ to $\overline{\mathbf{Q}}$. Let $V=\mathrm{Spa}(A_{V},A_{V}^{\circ})$ be the image of $U$ under the weight map (note that we shrink $U$, if necessary, so that $U$ is isomorphic to $V$) and let $\lambda_{V}:\mathbf{Z}_{p}^{\times}\rightarrow A_{V}^{\times}$ be the universal weight of $V$. Let $s\geq 1+s_{V}$ (cf. Remark A.2) and consider the étale cohomology $H^{1}_{\mathrm{\acute{e}t}}(\overline{Y},\mathscr{D}_{\lambda_{V}}^{s})$ (cf. Definition A.4). From the construction of $V({\mathbf{f}})$, one sees that it is a Hecke-equivariant direct summand of $H^{1}_{\mathrm{\acute{e}t}}(\overline{Y},\mathscr{D}_{\lambda_{V}}^{s})$ (and is independent to $s$). In particular, there exists a natural Hecke-equivariant projection
$\operatorname{Pr}_{{\mathbf{f}}}:H^{1}_{\mathrm{\acute{e}t}}(\overline{Y},\mathscr{D}_{\lambda_{V}}^{s})\rightarrow V({\mathbf{f}}).$
By [BC06, Thm. 1], $t_{{\mathbf{f}}}$ has maximal reducibility ideal. Then, it follows that the fiber of $V({\mathbf{f}})$ at $k=r$, which is exactly $V(f_{\beta})$, must be a non-split extension. The same applies to the other three modules, showing that all four of $V(f_{\beta}),V^{c}(f_{\beta}),V(f_{\beta})^{*}$ and $V^{c}(f_{\beta})^{*}$ are non-split extensions in the setting of Soulé’s theorem.
@@ 2.3. Crystalline periods
We begin by establishing some notation. Fix an embedding $\overline{\mathbf{Q}}\hookrightarrow\overline{\mathbf{Q}}_{p}$, that allow us to see $\operatorname{Gal}(\overline{\mathbf{Q}}_{p}/\mathbf{Q}_{p})$ as a subgroup of $\operatorname{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$. We write $\mathcal{R}$ for the Robba ring over $\mathbf{Q}_{p}$, $U=\mathrm{Spa}(A_{U},A_{U}^{\circ})$ be an affinoid open subset of $\mathcal{C}^{0}$, and let $\mathcal{R}_{U}=\mathcal{R}\mathop{\hat{\otimes}}_{\mathbf{Q}_{p}}A_{U}$. A $(\varphi,\Gamma)$-module is endowed with a Frobenius action (denoted by $\varphi$) and a Galois action, for which we take a topological generator $\gamma$ of $\operatorname{Gal}(\overline{\mathbf{Q}}_{p}/\mathbf{Q}_{p})$. As usual, let $t\in\mathcal{R}$ be the period for the cyclotomic character, so $\varphi(t)=pt$ and $\gamma(t)=\varepsilon(\gamma)t$. For $\delta\in A_{U}^{\times}$, let $\mathcal{R}_{U}(\delta)$ be the rank-1 $(\varphi,\Gamma)$-module over $\mathcal{R}_{U}$ generated by an element $e$ which is $\Gamma$-invariant and satisfies $\varphi(e)=\delta e$. [^1] We warn the reader that this is not the standard notation for rank-1 $(\varphi,\Gamma)$-modules. We use this notation to avoid introducing more technical terminology. Finally, when $D$ is a $(\varphi,\Gamma)$-module over $\mathcal{R}$ (or $\mathcal{R}_{U}$), write $\mathbf{D}_{\mathrm{cris}}(D)=D[1/t]^{\Gamma}$, with its filtration $\operatorname{Fil}^{n}\mathbf{D}_{\mathrm{cris}}(D)=(t^{n}D)^{\Gamma}$. Let $D({\mathbf{f}})^{*}=\mathbf{D}^{\dagger}_{\mathrm{rig}}(V({\mathbf{f}})^{*})$, and similarly $D^{c}({\mathbf{f}})^{*}$.
@@
is a finitely-generated $A_{U}$-module, by the main theorem of [KPX14]. In particular, it is free of rank 1, and there is a Zariski-dense set of $x\in A_{U}$ where the fiber is 1-dimensional. Let $\mathcal{F}^{+}D({\mathbf{f}})^{*}$ be the image of a generator of this map, and $\mathcal{F}^{-}$ the quotient, so that we have a short exact sequence
$0\to\mathcal{F}^{+}D({\mathbf{f}})^{*}\to D({\mathbf{f}})^{*}\to\mathcal{F}^{-}D({\mathbf{f}})^{*}\to 0.$
If $f_{\beta}$ is non-critical, the submodule $\mathcal{F}^{+}D(f_{\beta})^{*}$ is saturated, and $\mathcal{F}^{-}D({\mathbf{f}})^{*}$ is free of rank 1 over $\mathcal{R}_{U}$. Further, $\mathcal{F}^{\pm}D({\mathbf{f}})^{*}$ define a triangulation of $D({\mathbf{f}})^{*}$ over $U$.
@@
$b_{{\mathbf{f}}}^{+}:\mathbf{D}_{\mathrm{cris}}\left(\mathcal{F}^{+}D({\mathbf{f}})^{*}(-1-{\mathbf{k}})\right)\cong A_{U}.$
@@
Proposition 2.6.
There exists some scalar $c_{f_{\beta}}\in L^{\times}$ such that
$b_{f_{\beta}}^{+}=c_{f_{\beta}}\eta_{f_{\beta}}.$
Proof.
@@
and we have a duality between $V(f_{\beta})^{*}_{\mathrm{quo}}$ and $V^{c}(f_{\beta})_{\mathrm{quo}}$, where the subscript $\mathrm{quo}$ stands for the quotient in the corresponding filtrations, as defined in §2.2. This allows us to see $b_{f_{\beta}}^{+}$ as a basis of the space
$\mathbf{D}_{\mathrm{cris}}(V^{c}(f_{\beta})_{\mathrm{quo}})\cong H^{1}_{\mathrm{dR},\mathrm{c}}(Y,\mathscr{V}_{r})[\mathbb{T}=f_{\beta}]=H^{1}(X,\omega^{-r}(-\mathrm{cusps}))[\mathbb{T}=f_{\beta}],$
where the latter stands for the higher coherent cohomology of the compactification $X$ of $Y$ (cf. [LRS24, §A3]). Therefore, there exists some scalar $c_{f_{\beta}}\in L^{\times}$ such that
@@
although a priori we do not have any control over that constant. ∎
@@
$a_{{\mathbf{f}}}^{-}:\mathbf{D}_{\mathrm{cris}}\left(\mathcal{F}^{-}D({\mathbf{f}})^{*}\right)\cong A_{U}.$
In particular, if $g$ is a non-critical classical specialization of $\mathbf{f}$}, then $a_{{\mathbf{f}}}^{-}$ specializes to $d_{g}\eta_{g}$, where $d_{g}$ is some arbitrary non-zero constant.
Proposition 2.9.
There exists some scalar $d_{f_{\beta}}\in L^{\times}$ such that
$a_{f_{\beta}}^{-}=d_{f_{\beta}}\omega_{f_{\beta}}.$
Proof.
At $X=0$, we have an isomorphism
$\mathbf{D}_{\mathrm{cris}}(\mathcal{F}^{-}D(f_{\beta})^{*}(-1-r))\cong\mathbf{D}_{\mathrm{cris}}(V(f_{\beta})^{*}_{\mathrm{sub}}).$
Taking into account the dualities, and proceeding as before, we may claim that there exists a scalar $d_{f_{\beta}}\in L^{\times}$ such that $a_{f_{\beta}}^{-}=d_{f_{\beta}}\omega_{f_{\beta}}$. ∎
Remark 2.10.
The isomorphism we have used in the proof of the previous proposition is an isomorphism of the underlying $\varphi$-modules. Following the discussion after [LRS24, Prop. A8.1], there is a shift in the filtration degree by $r+1$.
@@ 3. Circular units
As in the introduction, let $\mathcal{H}_{\Gamma}$ be the distribution algebra of $\Gamma\cong\mathbf{Z}_{p}^{\times}$ (with $L$-coefficients); that is, $\mathcal{H}_{\Gamma}$ is the global section of the rigid analytic space associated with $\mathrm{Spf}\,\mathcal{O}_{L}[\![\Gamma]\!]$. For our further use, let ${\mathbf{j}}\colon\Gamma\hookrightarrow\mathcal{H}_{\Gamma}^{\times}$ be the universal character, which we regard as a character of the Galois group by composition with the cyclotomic character. The Kubota–Leopoldt $p$-adic $L$-function attached to $\chi$, denoted by $L_{p}(\chi)$, is an element of $\mathcal{H}_{\Gamma}$ which interpolates special values of the Dirichlet $L$-function of $\chi$. We refer [Das16, §3] for a more detailed treatment of the properties of this $p$-adic $L$-function with our current conventions, which slightly differ from [Col00].
@@
The element $L_{p}(\chi)$ belongs to the Iwasawa algebra $\mathcal{O}_{\chi,p}[\![\mathbf{Z}_{p}^{\times}]\!]$, which is a product of (p1)(p-1) subrings, indexed by $\mathbf{Z}/(p-1)\mathbf{Z}$, each of them isomorphic to $\mathcal{O}_{\chi,p}[\![T]\!]$.
Then, for any $t\in\mathbf{Z}/(p-1)\mathbf{Z}$ with $(-1)^{t}=\chi(-1)$, there exists a continuous function $L_{p,t}(\chi)\colon\mathbf{Z}_{p}\rightarrow\mathbf{C}_{p}$ such that, for all integers $k\geq 0$ with $k\equiv t\pmod{p-1}$, we have $L_{p,t}(\chi,k)=(1-\chi(p)p^{k-1})\cdot L(\chi,1-k)$. For the subrings corresponding to the sign condition $(-1)^{t}=-\chi(-1)$, we define the corresponding $L_{p,t}(\chi)$ using the interpolation property along the positive integers, thus having (p1)(p-1) different $p$-adic $L$-function which can be packaged in a single element $L_{p}(\chi)$ of $\mathcal{O}_{\chi,p}[\![\mathbf{Z}_{p}^{\times}]\!]$.
With these conventions, the Kubota–Leopoldt $p$-adic $L$-function satisfies a functional equation, which is discussed in [Das16, §3.2]. However, although the Kubota–Leopoldt $p$-adic $L$-function may be defined along any of the (p1)(p-1) discs of the weight space, regardless of any parity constraint, the cohomology classes can only be defined over half of them.
@@
Here, as usual, we identify $\chi$ with a character of the Galois group. It is characterized by the property that for any integer $j$, we recover $V(\chi)^{*}(j)=\mathcal{O}_{\chi,p}(\chi^{-1})(-j)$ by specializing at $j$ (viewed as a ring homomorphism $\mathcal{H}_{\Gamma}\to\mathbf{Q}_{p}$.